# Training Large Language Models to Reason In Parallel with Global Forking Tokens

### v1

@@ A.3.4 Ablation Study Training details (Removing High Quality Small Dataset)
For this ablation study, we choose Open-R1-Math220k default split, which has 93,000 math questions and $2\sim 4$ traces. Since Qwen2.5-Math-7B is a widely fine-tuned model using this dataset, we also choose it as our base model. We train for 3 epochs using 8 A100 GPUs, which took around 4 days. Our hyperparameters are mostly consistent with the recommended hyperparameters by Face (2025). We fine-tune both SSFT-OpenR1-93k-7B and SFT-OpenR1-93k-7B with a maximum length of 32768, learning rate of 4.0e054.0e-05 warmed up linearly for $3\%$ and decayed to 0 following cosine schedule, 8 gradient accumulation steps. For our SSFT method, we reserve $\mathrm{N}=4$ global forking tokens, and use the first 1000 tokens for matching. Again, only the completion loss is used for optimizing the model parameters.
@@ A.5 Alternative Implementation of training with multi-target sequences
Figure 11 shows an alternative implementation for training with multiple target parallel sequences. Instead of flattening multi-target data along the batch dimension, we can also concatenate parallel reasoning traces along the sequence dimension, modify the causal attention matrix, and position ids for positional embeddings. While easy to implement, this version extends the sequence length and cannot perform gradient accumulation along the batch dimension to maintain the same VRAM as training with a single reasoning trace. Since most fine-tuning already sets the micro-batch size to 1 due to the increased number of reasoning tokens already present in a single reasoning path, this implementation also cannot further shrink the micro-batch size to accommodate the extended context length. In addition, the code that uses the initial flash-attention implementation, which requires causal attention matrix, is not compatible with this setup. In terms of choosing the number of sequences in a global batch size, this implementation is also less flexible.

### v2

@@ A.3.4 Ablation Study Training details (Removing High Quality Small Dataset)
For this ablation study, we choose Open-R1-Math220k default split, which has 93,000 math questions and $2\sim 4$ traces. Since Qwen2.5-Math-7B is a widely fine-tuned model using this dataset, we also choose it as our base model. We train for 3 epochs using 8 A100 GPUs, which took around 4 days. Our hyperparameters are mostly consistent with the recommended hyperparameters by openr1. We fine-tune both SSFT-OpenR1-93k-7B and SFT-OpenR1-93k-7B with a maximum length of 32768, learning rate of 4.0e054.0e-05 warmed up linearly for $3\%$ and decayed to 0 following cosine schedule, 8 gradient accumulation steps. For our SSFT method, we reserve $\mathrm{N}=4$ global forking tokens, and use the first 1000 tokens for matching. Again, only the completion loss is used for optimizing the model parameters.
@@ A.5 Alternative Implementation of training with multi-target sequences
Figure 11 shows an alternative implementation for training with multiple target parallel sequences. Instead of flattening multi-target data along the batch dimension, we can also concatenate parallel reasoning traces along the sequence dimension, modify the causal attention matrix, and position ids for positional embeddings. While easy to implement, this version extends the sequence length and cannot perform gradient accumulation along the batch dimension to maintain the same VRAM as training with a single reasoning trace. Since most fine-tuning already sets the micro-batch size to 1 due to the increased number of reasoning tokens already present in a single reasoning path, this implementation also cannot further shrink the micro-batch size to accommodate the extended context length. In addition, the code that uses the initial flash-attention implementation, which requires causal attention matrix, is not compatible with this setup. This implementation is also less flexible in maintaining a consistent number of sequences per global batch.